# A demonstration of the  class file for SAGE Publications

### v1

@@ Abstract
Clothes manipulation, such as folding or hanging, is a critical capability for home service robots. Despite recent advances, most existing methods remain limited to specific tasks and clothes types, due to the complex, high-dimensional geometry of clothes. This paper presents CLothes mAnipulation with Semantic keyPoints (CLASP), which aims at general-purpose clothes manipulation over different clothes types, T-shirts, shorts, skirts, long dresses, … , as well as different tasks, folding, flattening, hanging, … . The core idea of CLASP is semantic keypoints—e.g., “left sleeve”, “right shoulder”, etc.—a sparse spatial-semantic representation that is salient for both perception and action. Semantic keypoints of clothes can be reliably extracted from RGB-D images and provide an effective intermediate representation of clothes manipulation policies. CLASP uses semantic keypoints to bridge high-level task planning and low-level action execution. At the high level, it exploits vision language models (VLMs) to predict task plans over the semantic keypoints. At the low level, it executes the plans with the help of a simple pre-built manipulation skill library. Extensive simulation experiments show that CLASP outperforms state-of-the-art baseline methods on multiple tasks across diverse clothes types, demonstrating strong performance and generalization. Further experiments with a Franka dual-arm system on four distinct tasks—folding, flattening, hanging, and placing—confirm CLASP’s performance on a real robot. The CLASP source code is available at https://github.com/dengyh16code/CLASP.
@@ 1 Introduction
People have long anticipated that an intelligent home service robot would take care of their laundry chores, like folding a T-shirt for storage or hanging a skirt on a hanger. Toward this goal, we need a general-purpose robot capable of manipulating a wide variety of clothes in many different ways. Despite recent significant advances in clothes manipulation skills like folding and flattening (Moletta et al., 2023; Canberk et al., 2023), these methods are tailored to specific clothes types and tasks.
How can we build a robot for general-purpose clothes manipulation? Clothes are deformable objects with a high-dimensional state space, and different types of clothes (e.g., T-shirts and skirts) exhibit distinct geometric structures. This complexity makes it essential to develop an effective and general state representation for general-purpose manipulation. Although physics-based representations, such as particles and meshes (Bender et al., 2015; Chen et al., 2022), can represent a wide range of clothes, they are complex and challenging to estimate accurately, especially under deformation and occlusion. Keypoint-based representation provides a more succinct representation, which uses dense keypoints to capture the topological structure of clothes (Gustavsson et al., 2022; Lips et al., 2024). However, these dense keypoints are also challenging to track and estimate under deformation and occlusion. Additionally, these keypoints are tailored to specific clothes items and do not generalize well across different instances. These limitations arise because previous keypoint representations focus only on geometric information and ignore semantic information, which is crucial given that clothes are manually designed and inherently shaped for humans.
In this paper, we present semantic keypoints as a general spatial-semantic representation of clothes. Compared with previous keypoint representations, semantic keypoints carry explicit semantic meaning and can be intuitively described using natural language. As a result, semantic keypoints offer a more sparse and succinct representation and focus on distinct structural features of clothes like sleeves and shoulders. These structural features are salient for both perception and action. In the context of perception, semantic keypoints are relatively easy to extract from observations and remain consistent across diverse clothes instances of the same type. This consistency is the foundation of generalization. In the context of action, semantic keypoints not only capture the geometric structure of clothes but also identify where clothes are frequently manipulated. As a result, semantic keypoints provide an effective abstraction of the action space for task planning, bridging the gap between perception and manipulation.
@@
For general-purpose clothes manipulation, we integrate semantic keypoints with foundation models and develop a CLothes mAnipulation method with Semantic keyPoints (CLASP). As Figure 1 shows, CLASP can be deployed on a wide range of clothes types and manipulation tasks. CLASP first leverages foundation models for open-category semantic keypoint extraction. Semantic keypoints are the inherent features of clothes, which can be effectively captured in the commonsense knowledge of foundation models. Specifically, we propose a two-stage pipeline for semantic keypoint extraction. The first stage leverages a vision language model (VLM) for semantic understanding of clothes and autonomously discovers semantic keypoints on a prototype image for each clothes category. The second stage aims to match semantic keypoints on the prototype image to novel clothes, where vision foundation models ensure spatial precision under deformation and occlusion.
To manipulate clothes using semantic keypoints, we first build a skill library. We leverage the commonsense knowledge of a large language model (LLM) to identify general basic skills for clothes manipulation, then implement them as policies parameterized by contact point positions. Given the skill library, CLASP leverages a VLM-powered task planner to decompose the free-form natural language instruction into sequential subtasks conditioned on the observation image and semantic keypoints. Each subtask is described by a basic skill selected from the skill library and contact points like grasp(“left sleeve”). In task planning, semantic keypoints define where clothes can be manipulated and serve as contact point candidates. After obtaining the task plan, CLASP verifies whether the plan is executable. If the task plan fails verification, the failed subtask prompts the VLM to replan; otherwise, CLASP executes the subtasks sequentially. For each subtask, CLASP invokes a policy from the skill library to generate waypoints based on contact point positions and applies motion planning to produce an executable trajectory. After executing each subtask, CLASP updates the observation and determines whether to replan. This closed-loop pipeline will be repeated until the task is completed.
To evaluate CLASP, we extend the SoftGym (Lin et al., 2021) benchmark and conduct simulation experiments across four clothes categories and four representative tasks in various scenarios. CLASP outperforms baseline methods like flingbot (Ha and Song, 2022) and fabricflownet (Weng et al., 2022) in clothes manipulation. Real-world experiments on 15 various clothes show that CLASP can be easily deployed on a real robot, performing well across a wide variety of clothes and tasks. These results demonstrate that semantic keypoints are an effective representation of clothes manipulation, enabling general-purpose clothes manipulation with foundation models.
This paper is an extension of our previous work presented at IEEE International Conference on Robotics and Automation (ICRA), titled “General-purpose Clothes Manipulation with Semantic Keypoints” (Deng and Hsu, 2025). We expand upon the conference version in three primary aspects: (i) Open-category semantic keypoint extraction: we improve the semantic keypoint extraction method by leveraging foundation models, eliminating the need for training on labeled data. This enhancement significantly broadens the applicability and generalization of our method to open-category clothes. (ii) Closed-loop task planning: we improve the previous open-loop system with plan verification, dynamic replanning from new observations, and task completion checks—greatly improving robustness to unexpected state changes, especially with initially crumpled clothes. (iii) Expanded Real-world Evaluations: we extend the scope of real-world experiments to a diverse range, including T-shirts, sweaters, shorts, pants, skirts, long dresses, and more. Additionally, we conducted comprehensive quantitative real-world experiments and analysis to validate the effectiveness and generalization of our method.
@@ 2.2 Deformable Object Manipulation
Existing deformable manipulation approaches can be categorized into model-based and model-free approaches. Model-based approaches use a dynamic model to predict configurations of deformable objects under a certain action and then select the appropriate action (Cusumano-Towner et al., 2011; Miller et al., 2012; Yu et al., 2023). The key challenge lies in the dynamic modeling of deformable objects. To tackle this, two main techniques are used: physics-based models (Li et al., 2018; Chen et al., 2024), which rely on physical law, and data-driven models (Tanaka et al., 2018; Lin et al., 2022), which are learned from interaction data. Typical physics-based models include particle-based models (Longhini et al., 2024b; Zhang et al., 2024b) and mesh-based models (Huang et al., 2022). Recently, Material Point Method (MPM) (Jiang et al., 2016) has emerged as a new physics-based modeling method, which combines particles with grid-based forces for more accurate dynamic modeling (Gu et al., 2023). However, physics-based modes often struggle with heavy parameter tuning (Zheng et al., 2024). Data-driven dynamics models are learned from interaction data, typically formulated as a supervised learning problem (Hu et al., 2018; Tanaka et al., 2018; Yan et al., 2021). Compared to physics-based models, data-driven models support more flexible state space, especially via latent representation (Lippi et al., 2020; Li et al., 2024). However, data-driven models rely on expensive robot interaction data and often fail to generalize to different objects and tasks. Overall, the generalization of model-based approaches highly depends on accurate dynamic modeling and calibration for unseen objects and tasks (Longhini et al., 2024a), which remains an open problem due to the complex, nonlinear dynamics of deformable objects.
@@
Although model-free approaches have made significant progress in a wide range of tasks (Ganapathi et al., 2021; Xu et al., 2022; Bahety et al., 2023), the generalization remains a significant limitation. Most of the previous methods aim to learn task-specific policies (Salhotra et al., 2022; Weng et al., 2022; Gao et al., 2023). Although some goal-conditioned methods use goal images for multi-task learning (Seita et al., 2021; Chi et al., 2022), they often struggle to generalize to unseen goals. To enhance generalization, collecting more data (Xue et al., 2023) and domain randomization (Hietala et al., 2022; Wang et al., 2023) have been explored. However, model-free approaches still struggle to generalize to unseen tasks or object categories. In this paper, we aim to tackle general-purpose cloth manipulation across diverse clothes categories and task types. To this end, we propose CLASP, a framework that integrates semantic keypoint representations with foundation models. CLASP enables generalization to a wide range of clothes and manipulation tasks.
Figure 2: CLASP Overview. Given an RGB-D observation, CLASP extracts semantic keypoints as the state. These keypoints, along with the RGB image and task instruction, are fed to a VLM to generate a task plan. Once the plan is verified, the subtasks are executed sequentially. After each subtask, CLASP updates the observation and determines whether to replan. This process repeats until the overall task is completed.
@@ 2.3 Language-Conditioned Manipulation
Language-conditioned manipulation is an emerging field aimed at enabling seamless interaction between humans and robots by teaching robots to understand and execute instructions specified in natural language (Shridhar et al., 2022; Mees et al., 2022; Zhou et al., 2023b). The recent emergence of large foundation models, such as large language models (LLMs) and vision-language models (VLMs), has significantly advanced language-conditioned manipulation tasks. These models provide commonsense knowledge for reasoning (Singh et al., 2023; Du et al., 2024) and perception (Huang et al., 2023; Gao et al., 2024), significantly enhancing generalization capabilities.
Beyond specifying tasks, natural language also serves as a powerful abstraction of affordance (Mees et al., 2023), capturing the functions of an object, i.e., how it can be manipulated. For task-oriented grasping (Liu et al., 2020), GraspGPT (Tang et al., 2023) utilizes object class descriptions generated by a large language model (LLM) to identify optimal grasping regions. Similarly,  Ren et al. (2023) provide natural language descriptions of tools to accelerate learning language-conditioned tool manipulation policies. RoboPoint fine-tunes a VLM for spatial affordance prediction conditioned on natural language (Yuan et al., 2024).
@@ 3 Overview
In this paper, we introduce CLASP (CLothes mAnipulation method with Semantic keyPoints), a general-purpose clothes manipulation method. The core idea of CLASP is to use semantic keypoints as a general state representation for clothes. Each semantic keypoint includes: (i) a language description (e.g., “left sleeve”), providing semantic context for high-level task planning, and (ii) a keypoint position, offering geometric information to guide low-level action generation.
To leverage semantic keypoints in clothes manipulation, we first build a basic skill library. Instead of manually designing skills, we prompt an LLM to discover basic clothes manipulation skills. The LLM’s commonsense knowledge ensures these skills are general and sufficient. We then implement the discovered skills as policies parameterized by contact point positions, to form the skill library.
Fig. 2 illustrates the overall framework of CLASP. Given RGB-D observations of clothes and a language instruction, CLASP generates sequential manipulation actions to complete the task specified by the instruction. CLASP first extracts semantic keypoints through foundation models. After extracting semantic keypoints, the RGB image marked with semantic keypoints and the language instruction are fed into a VLM for task planning. The VLM determines whether the task is complete and infers a task plan, where each subtask consists of a basic skill and contact points. After obtaining the task plan, CLASP verifies whether the plan is executable. The verified task plan will be executed sequentially by invoking policies from the skill library. After each subtask, CLASP updates the observation and decides whether to replan. The closed-loop pipeline repeats until the task is completed.
In the following subsections, we will introduce how we extract effective semantic keypoints of clothes in Sec. 4. Then, to enable general-purpose clothes manipulation with semantic keypoints, we present the construction of the skill library in Sec. 5.1, followed by the details of task planning and execution in Sec 5.2.
@@ 4 Semantic Keypoints
Although previous work has explored using keypoints for clothes manipulation, most methods rely on manually annotated keypoints (Mo et al., 2022; Lips et al., 2024). Since clothes are highly deformable, collecting sufficient data that captures clothes’ various configurations is time-consuming and labor-intensive. To mitigate data scarcity, augmenting 3D synthetic data in the simulation environment has been explored (Deng and Hsu, 2025). However, this approach is difficult to scale to open-category clothes due to the limited availability of 3D clothes models. Recently, researchers have explored zero-shot keypoint extraction by leveraging powerful and generalizable feature maps from vision foundation models (Huang et al., 2024; Di Palo and Johns, 2024; Fang et al., 2024b). However, these methods can not ensure that the resulting keypoints have explicit semantic meaning and can be described using natural language. In addition, using the vision foundation model for consistent keypoint extraction on clothes remains challenging, especially under self-occlusion and deformation.
Given that semantic keypoints encode both semantic and spatial information, two key technical challenges must be addressed in keypoint extraction: understanding the semantics of open-category clothes and maintaining the spatial precision of keypoints under deformation and occlusion. To address them, we propose a two-stage pipeline for keypoint extraction. The first stage (Fig. 3) focuses on semantic understanding, aiming to autonomously discover semantic keypoints on prototype images and describe them using natural language. Specifically, we leverage geometric prior and the vision language model’s commonsense knowledge to develop an autonomous semantic keypoint discovery pipeline. For each clothes category, we provide a prototype image free from occlusion and deformation, which makes semantic understanding feasible and also provides full observation as a reference in the second stage. The second stage (Fig. 5) focuses on spatial precision, aiming to match semantic keypoints on the prototype image to novel clothes, which may involve occlusions and deformations. Specifically, we utilize a coarse-to-fine matching strategy to propose semantic keypoints on novel clothes precisely.
Figure 3: Semantic Keypoint Detection. We design an autonomous pipeline to discover semantic keypoints on a prototype image without occlusion and deformation for each clothes type.
Figure 4: Prototype Image Retrieval. Given an RGB observation image of novel clothes, we retrieve the most relevant prototype image by computing Best Buddies Similarity (BBS) using DINOv2 features.
Figure 5: Semantic Keypoint Matching. Given the prototype image and discovered semantic keypoints, we matched the discovered semantic keypoints to novel clothes with VLM and vision foundation models.
4.1 Semantic Keypoint Discovery

### v2

# CLASP: General-Purpose Clothes Manipulation with Semantic Keypoints

@@ Abstract
Clothes manipulation, such as folding or hanging, is a critical capability for home service robots. Despite recent advances, most existing methods remain limited to specific clothes types and tasks, due to the complex, high-dimensional geometry of clothes. This paper presents CLothes mAnipulation with Semantic keyPoints (CLASP), which aims at general-purpose clothes manipulation over diverse clothes types, T-shirts, shorts, skirts, long dresses, … , as well as different tasks, folding, flattening, hanging, … . The core idea of CLASP is semantic keypoints—e.g., “left sleeve” and “right shoulder”—a sparse spatial-semantic representation, salient for both perception and action. Semantic keypoints of clothes can be reliably extracted from RGB-D images and provide an effective representation for a wide range of clothes manipulation policies. CLASP uses semantic keypoints as an intermediate representation to connect high-level task planning and low-level action execution. At the high level, it exploits vision language models (VLMs) to predict task plans over the semantic keypoints. At the low level, it executes the plans with the help of a set of pre-built manipulation skills conditioned on the keypoints. Extensive simulation experiments show that CLASP outperforms state-of-the-art baseline methods on multiple tasks across diverse clothes types, demonstrating strong performance and generalization. Further experiments with a Franka dual-arm system on four distinct tasks—folding, flattening, hanging, and placing—confirm CLASP’s performance on real-life clothes manipulation. The CLASP source code is available online at https://github.com/dengyh16code/CLASP.
@@ 1 Introduction
People have long anticipated intelligent service robots performing home laundry chores: fold or hang shirts, pants, gowns … . While there has been significant progress in robot clothes manipulation in recent years (Avigal et al., 2022; Huang et al., 2022; Canberk et al., 2023; Longhini et al., 2024b), existing methods are usually tailored to specific clothes and tasks. General-purpose clothes manipulation—a single robot system performing different manipulation tasks over diverse clothes types—remains an important open challenge.
A core issue here is representation. Clothes are deformable objects with high-dimensional state space and complex geometry that varies significantly across categories, e.g., shirts and gowns. How do we represent clothes? Dense physics-based representations, such as particles (Bender et al., 2015; Chen et al., 2022), can model diverse clothes, but the fine details required are difficult to estimate from visual observations, because of clothes deformation and self-occlusion (Yin et al., 2021). Sparse keypoints (Gustavsson et al., 2022; Lips et al., 2024) are more robust to deformation, but generalize poorly over different clothes (Longhini et al., 2024a). Both approaches focus exclusively on geometry, while ignoring semantics. Semantics is crucial for clothes representation, as clothes are man-made objects designed specifically for humans.
We introduce semantic keypoints (Figure 1), e.g., “left sleeve” and “right shoulder”, as a spatial–semantic representation that connects perception and action for general-purpose clothes manipulation. Semantic keypoints provide significant advantages. First, they are salient for both perception and action. They can be extracted reliably with RGB-D sensors under moderate deformation and occlusion. They also identify commonly manipulated regions on clothes and highlight their affordance, providing an abstraction to guide action planning and execution. Second, semantic keypoint descriptors naturally emerge in human language due to their salience. Finally, semantic keypoints are grounded to structural features consistent within the same clothes category, enabling effective generalization.
@@
Semantic keypoint extraction poses two main challenges: (i) identifying semantically meaningful keypoints for open-category clothes, and (ii) achieving accurate spatial localization of keypoints under deformation and occlusion. To decouple the semantic and spatial challenges, we propose a two-stage method. First, we use a vision-language model (VLM) for semantic understanding of individual clothes categories. For each category, we determine the semantic keypoints on a prototype clothes instance lying flat. Next, we match these keypoints spatially to novel input clothes instances under possible deformation and occlusion, using vision foundation models.
Given a natural language instruction and RGB-D observations, CLothes mAnipulation with Semantic keyPoints (CLASP) performs various manipulation tasks over a wide range of clothes types. To tackle the challenge of general-purpose clothes manipulation, CLASP leverages semantic keypoints to decompose the manipulation task into a two-level hierarchy (Figure 2). At the high level, CLASP uses a VLM to transform the input language instruction into a task plan over the semantic keypoints. For example, grasp the left sleeve, move to the center, … . It verifies the generated task plan for geometric feasibility through motion planning and decides whether to replan. At the low level, CLASP executes each subtask in the high-level plan, using a set of pre-built manipulation skills conditioned on the keypoints. After executing each subtask, CLASP closes the loop by updating the visual observation and replanning if needed. The process continues until the task is successfully completed. To evaluate CLASP, we conducted experiments both in simulation and on a dual-arm robot system. For simulation experiments, we extended the SoftGym (Lin et al., 2021) benchmark and experimented with four clothes categories and four representative tasks under various scenarios. CLASP outperforms well-established baselines in clothes manipulation, including FlingBot (Ha and Song, 2022) and FabricFlownet (Weng et al., 2022). For real-robot experiments, we implemented CLASP on Franka robots for dual-arm manipulation on 15 clothes of various types. The results confirm the performance of CLASP and its generality. They demonstrate that semantic keypoints provide an effective representation of clothes manipulation, enabling general-purpose clothes manipulation with foundation models.
This work extends our earlier paper (Deng and Hsu, 2025) in three main aspects:
  • Semantic keypoint extraction for open-category clothes. We improved semantic keypoint extraction by leveraging foundation models, eliminating the need for labeled training data (Section 4). This significantly enhances the generalization performance of CLASP on open-category clothes.
  • Closed-loop planning. CLASP now closes the loop at both the task planning and the subtask execution levels (Section 5.2). Closed-loop planning substantially improves overall system performance (Sec. 6.2), especially with crumpled clothes.
  • Expanded real-world evaluation. We added diverse clothes, such as T-shirts, sweaters, shorts, pants, skirts, and dresses, and conducted comprehensive real-robot experiments to validate the effectiveness and generalization of CLASP (Section 6.4).
The rest of this paper is organized as follows. The related work is reviewed in Section 2. Section 3 introduces the overall CLASP framework. Section 4 presents semantic keypoints extraction. Section 5 details clothes manipulation guided by semantic keypoints. The experimental setup and results are provided in Section 6. Section 7 discusses limitations and future directions. Finally, Section 8 concludes the paper. Additional implementation and experimental details are provided in the appendix.
@@ 2.2 Deformable Object Manipulation
Existing deformable manipulation approaches can be categorized into model-based and model-free approaches. Model-based approaches use a dynamic model to predict configurations of deformable objects under a certain action and then select the appropriate action (Cusumano-Towner et al., 2011; Miller et al., 2012; Yu et al., 2023). The key challenge lies in the dynamic modeling of deformable objects. To tackle this, two main techniques are used: physics-based models (Li et al., 2018; Chen et al., 2024), which rely on physical law, and data-driven models (Tanaka et al., 2018; Lin et al., 2022), which are learned from interaction data. Typical physics-based models include particle-based models (Longhini et al., 2024b; Zhang et al., 2024b) and mesh-based models (Huang et al., 2022). Recently, Material Point Method (MPM) (Jiang et al., 2016) has emerged as a new physics-based modeling method, which combines particles with grid-based forces for more accurate dynamic modeling (Gu et al., 2023). However, physics-based modes often struggle with heavy parameter tuning (Zheng et al., 2024). Data-driven dynamics models are learned from interaction data, typically formulated as a supervised learning problem (Hu et al., 2018; Tanaka et al., 2018; Yan et al., 2021). Compared to physics-based models, data-driven models support a more flexible state space, especially via latent representation (Lippi et al., 2020; Li et al., 2024). However, data-driven models rely on expensive robot interaction data and often fail to generalize to different objects and tasks. Overall, the generalization of model-based approaches highly depends on accurate dynamic modeling and calibration for unseen objects and tasks (Longhini et al., 2024a), which remains an open problem due to the complex, nonlinear dynamics of deformable objects.
@@
Although model-free approaches have made significant progress in a wide range of tasks (Ganapathi et al., 2021; Xu et al., 2022; Bahety et al., 2023), the generalization remains a significant limitation. Most of the previous methods aim to learn task-specific policies (Salhotra et al., 2022; Weng et al., 2022; Gao et al., 2023). Although some goal-conditioned methods use goal images for multi-task learning (Seita et al., 2021; Chi et al., 2022), they often struggle to generalize to unseen goals. To enhance generalization, collecting more data (Xue et al., 2023) and domain randomization (Hietala et al., 2022; Wang et al., 2023) have been explored. However, model-free approaches still struggle to generalize to unseen tasks or object categories. In this paper, we aim to tackle general-purpose clothes manipulation across diverse clothes categories and task types. To this end, we propose CLASP, a framework that integrates semantic keypoint representations with foundation models. CLASP enables generalization to a wide range of clothes and manipulation tasks.
Figure 2: CLASP Overview. Given an RGB-D observation, CLASP extracts semantic keypoints. These keypoints, along with the RGB image and task instruction, are fed to a VLM to generate a task plan. Once the plan is verified for feasibility through motion planning, the subtasks are executed sequentially. After each subtask execution, CLASP updates the observation and decides whether to replan. This process repeats until the overall task is completed.
@@ 2.3 Language-Conditioned Manipulation
Language-conditioned manipulation is an emerging field that aims to enable seamless interaction between humans and robots by teaching robots to understand and execute instructions specified in natural language (Shridhar et al., 2022; Mees et al., 2022; Zhou et al., 2023b). The recent emergence of large foundation models that are pretrained on internet-scale data, such as LLMs and VLMs, has significantly advanced language-conditioned manipulation tasks. These models provide commonsense knowledge for reasoning (Singh et al., 2023; Du et al., 2024) and perception (Huang et al., 2023; Gao et al., 2024), significantly enhancing generalization capabilities. More recently, vision–language–action models (VLAs) (Kim et al., 2024; Black et al., 2025; Pertsch et al., 2025) fine-tune VLMs for robot control, aiming for generalist policies but relying on large-scale robotic data.
Beyond specifying tasks, natural language also serves as a powerful abstraction of affordance (Mees et al., 2023), capturing the functions of an object, i.e., how it can be manipulated. For task-oriented grasping (Liu et al., 2020), GraspGPT (Tang et al., 2023) utilizes object class descriptions generated by a large language model (LLM) to identify optimal grasping regions. Similarly,  Ren et al. (2023) provide natural language descriptions of tools to accelerate learning language-conditioned tool manipulation policies. RoboPoint finetunes a VLM for spatial affordance prediction conditioned on natural language (Yuan et al., 2024). MimicFunc (Tang et al., 2025) leverages functional cues to identify keypoints for tool manipulation.
@@ 3 Overview
For general-purpose clothes manipulation, CLASP leverages semantic keypoints to decompose the task into a two-level hierarchy. Figure 2 illustrates the overall framework. Given RGB-D observations of clothes and a language instruction, CLASP generates a sequence of manipulation actions to complete the specified task.
First, CLASP extracts semantic keypoints using foundation models. Each extracted keypoint is represented by a 3D position and an associated language descriptor. These keypoints, together with the RGB image and task instruction, are then fed into a vision-language model (VLM) to generate a task plan over semantic keypoints (e.g., grasp(left sleeve), moveto(center), …). Each subtask in the plan is grounded by selecting the appropriate skill from a pre-built skill library. The plan is then verified for feasibility through motion planning, which generates the corresponding robot trajectories for execution. After each subtask, CLASP updates its observation and determines whether replanning is required. This closed-loop process continues until the task is successfully completed.
In the following sections, we will introduce how we extract effective semantic keypoints of open-category clothes in Sec. 4. Then, to enable general-purpose clothes manipulation with semantic keypoints, we present the construction of the skill library in Sec. 5.1, followed by the details of task planning and execution in Sec 5.2.
Figure 3: Semantic keypoint discovery. A fully-automated pipeline discovers semantic keypoints on a prototype image for each clothes type.
@@ 4 Semantic Keypoints
The core idea of CLASP is to use semantic keypoints as a general state representation for clothes. Each semantic keypoint includes: (i) a language description (e.g., “left sleeve”), providing semantic context for high-level task planning, and (ii) a keypoint position, offering geometric information to guide low-level action generation. Although previous work has explored using keypoints for clothes manipulation, most methods rely on manually annotated keypoints (Mo et al., 2022; Lips et al., 2024). Since clothes are highly deformable, collecting sufficient data that captures clothes’ various configurations is time-consuming and labor-intensive. To mitigate data scarcity, augmenting 3D synthetic data in the simulation environment has been explored (Deng and Hsu, 2025). However, this approach is difficult to scale to open-category clothes due to the limited availability of 3D clothes models. Recently, researchers have explored zero-shot keypoint extraction by leveraging powerful and generalizable feature maps from vision foundation models (Huang et al., 2024; Di Palo and Johns, 2024; Fang et al., 2024b). However, these methods can not ensure that the resulting keypoints have explicit semantic meaning and can be described using natural language. In addition, using the vision foundation model for consistent keypoint extraction on clothes remains challenging, especially under self-occlusion and deformation.
Given that semantic keypoints encode both semantic and spatial information, two key technical challenges must be addressed in keypoint extraction: (1) understanding the semantics of open-category clothes and (2) maintaining the spatial precision of keypoints under deformation and occlusion. To address them, we propose a two-stage pipeline for keypoint extraction. The first stage (Figure 3) focuses on semantic understanding, aiming to autonomously discover semantic keypoints on prototype images and describe them using natural language. Specifically, we leverage geometric prior and the vision language model’s commonsense knowledge to develop an autonomous semantic keypoint discovery pipeline. For each clothes category, we provide a prototype image free from occlusion and deformation, which makes semantic understanding feasible and also provides full observation as a reference in the second stage. The second stage (Figure 4) focuses on spatial precision, aiming to match semantic keypoints on the prototype image to novel clothes, which may involve occlusions and deformations. Specifically, we utilize a coarse-to-fine matching strategy to propose semantic keypoints on novel clothes precisely.
Figure 4: Semantic keypoint matching. Given an image of novel clothes, we first retrieve the most relevant prototype. Each semantic keypoint on the prototype is then matched to the novel clothes through a coarse-to-fine pipeline. Specifically, a VLM is employed for coarse region matching, while Stable Diffusion (SD) and DINOv2 are utilized for fine-grained keypoint matching.
4.1 Keypoint Discovery